# Field of values analysis that includes zero for preconditioned nonsymmetric saddle-point systems

### v1

@@ 3 Preconditioned Saddle-Point Matrices with Zero in the Field of Values
where $F\in\mathbb{R}^{n\times n}$ is nonsingular and $B\in\mathbb{R}^{m\times n}$ is full row rank. We assume that $F$ is nonsymmetric and positive real (or positive definite), namely, that $u^{T}Fu>0$ for all $0\neq u\in{\mathbb{R}}^{n}$.
@@
Let $\alpha$ be a constant independent of the matrix dimensions, such that
$\|N\|_{H_{1},H_{1}^{-1}}\leq\alpha.$ (9)
@@
The result can be readily deduced by using the standard property of norms
$\|F^{-1}\|_{H_{1}^{-1},H_{1}}^{-1}$ $=\min_{v\in\mathbb{R}^{n}\backslash\{0\}}\max_{w\in\mathbb{R}^{n}\backslash\{0\}}\frac{w^{T}Fv}{\|v\|_{H_{1}}\|w\|_{H_{1}}}$
$\geq\min_{v\in\mathbb{R}^{n}\backslash\{0\}}\frac{v^{T}Fv}{\|v\|_{H_{1}}^{2}}\geq 1,$
which is stated in result (ii) of Lemma 2.2 (see also [17, Lemma 1]). ∎
@@
$\qquad\|B\|_{H_{1},H_{2}^{-1}}\leq C_{1},\qquad\min_{x}\frac{\|B^{T}x\|_{H_{1}^{-1}}}{\|x\|_{H_{2}}}\geq C_{2},$ (10)
where $C_{1}$ and $C_{2}$ are independent of $\alpha$ and the dimensions of $K$.
@@
$\|S^{-1}\|_{H_{2}^{-1},H_{2}}^{-1}$ $=\min_{v\in\mathbb{R}^{m}\backslash\{0\}}\max_{w\in\mathbb{R}^{m}\backslash\{0\}}\frac{w^{T}BF^{-1}B^{T}v}{\|v\|_{H_{2}}\|w\|_{H_{2}}}$
$\geq\min_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{v^{T}BF^{-1}B^{T}v}{\|v\|_{H_{2}}^{2}}$
$\geq\min_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{v^{T}BF^{-1}B^{T}v}{v^{T}BH_{1}^{-1}B^{T}v}\min_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{\|B^{T}v\|_{H_{1}^{-1}}^{2}}{\|v\|_{H_{2}}^{2}}$
$\geq C_{2}^{2}\min_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{v^{T}BF^{-1}B^{T}v}{v^{T}BH_{1}^{-1}B^{T}v}.$
Using [17, Lemma 8] and lemma 3.3, we have
$\min_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{v^{T}BF^{-1}B^{T}v}{v^{T}BH_{1}^{-1}B^{T}v}$ $\geq\min_{y\in\mathbb{R}^{n}\backslash\{0\}}\frac{y^{T}(I+H_{1}^{-1/2}NH_{1}^{-1/2})^{-1}y}{y^{T}y}$
@@
$\geq\frac{1}{(1+\alpha)^{2}},$
which completes the proof. ∎
@@
For a matrix $T$ and a scalar $\tau$, we say that $\|T\|\lesssim\tau$ if $\|T\|\leq C\tau$, where $C$ is a constant independent of the dimensions of $T$.
@@ 3.1 Block-Triangular Preconditioners
It is well known that there are some differences in the use of left and right preconditioners. For example, in flexible GMRES it is necessary to use right preconditioning. The correct norm considered in GMRES for finite element discretizations should typically be $\|\cdot\|_{H^{-1}}$ [1].
@@
$(M_{U}^{-1}K)^{-1}=\begin{bmatrix}I-F^{-1}B^{T}S^{-1}B&F^{-1}B^{T}S^{-1}H_{2}\\ S^{-1}B&I-S^{-1}H_{2}\end{bmatrix}.$
We now need to establish conditions (2a)–(2d) of Lemma 2.4.
@@
The proof is obtained by putting together the bounds of Lemmas 3.3, 3.4, and 3.6.
@@
$\leq(1+2C_{1}^{2}{(1+{\alpha})^{2}/C_{2}^{2}})+({(1+{\alpha})^{2}/C_{2}^{2}}C_{1})+1+{(1+{\alpha})^{2}/C_{2}^{2}}$
$\lesssim 1.$
@@
When $\alpha<\frac{1}{2}$, we have
$\left\|H(M_{U}^{-1}K)-(M_{U}^{-1}K)^{T}H\right\|_{H,H^{-1}}\lesssim{\alpha}.$ (13)
@@
When ${\alpha}<\frac{1}{2}$ we have
$\|B_{12}\|_{2}\leq C_{1}\frac{\|N\|_{H_{1},H_{1}^{-1}}}{1-\|N\|_{H_{1},H_{1}^{-1}}}\leq C_{1}{\alpha}/(1-{\alpha})\leq 2C_{1}\alpha\lesssim\alpha$ (15)
@@
$\leq 2C_{1}^{2}{\alpha}$
$\lesssim{\alpha}.$
Substituting the above inequalities into (14), we obtain (13), as required.
The results of Lemmas 3.83.10 along with the assumption that ${\alpha}$ is sufficiently small establish the scalability of the iterations.
@@
Given a saddle-point system with matrix $K$ defined in (5), where $F\in\mathbb{R}^{n\times n}$ is positive real and $B\in\mathbb{R}^{m\times n}$ is full row rank, let $H_{1}$ and $N$ be the symmetric and skew-symmetric parts, respectively, of $F$, as in (8). Let $H_{2}$ be a symmetric positive definite matrix, such that the three conditions in (9)–(10) are satisfied. Finally, let $H$ be the block-diagonal matrix defined in (7). Then, for $\alpha$ sufficiently small, GMRES with the left preconditioner $M_{U}$ under the $H$-norm will converge in a fixed number of iterations independently of the dimensions of $K$.
@@
Lemmas 3.83.10 establish conditions (2a)–(2c). Trivially, by Lemmas 3.9 and 3.10, (2d) holds when ${\alpha}$ is sufficiently small.
We now consider the right preconditioner $M_{L}$ defined in (12). The analysis is very similar to the left preconditioner case. The details are omitted and we present a theorem analogous to Theorem 3.11.
@@
Given a saddle-point system with matrix $K$ defined in (5), where $F\in\mathbb{R}^{n\times n}$ is positive real and $B\in\mathbb{R}^{m\times n}$ is full row rank, let $H_{1}$ and $N$ be the symmetric and skew-symmetric parts, respectively, of $F$, as in (8). Let $H_{2}$ be a symmetric positive definite matrix, such that the three conditions in (9)–(10) are satisfied. Finally, let $H$ be the block-diagonal matrix defined in (7). Then, for $\alpha$ sufficiently small, GMRES with the right preconditioner $M_{L}$ under the $H^{-1}$-norm will converge in a fixed number of iterations independently of the dimensions of $K$.
@@ 3.2 A Block-Diagonal Preconditioner
is interesting in the context of this work, because contrary to block-triangular preconditioners, where one might select either an upper block-triangular preconditioner or a lower block-triangular preconditioner along with left or right preconditioning to avoid a situation of having zero in the field of values, here it is immediate that the field of values contains zero regardless of any such choices made. There is no practical difference between left and right preconditioning here, and we proceed with left preconditioning below. The preconditioned matrix is
@@
The analysis is essentially identical to the analysis in Section 3.1.
@@
When $\alpha<\frac{1}{2}$, we have
$\left\|H(M_{2}^{-1}K)-(M_{2}^{-1}K)^{T}H\right\|_{H,H^{-1}}\lesssim\alpha.$
@@
Given a saddle-point system with matrix $K$ defined in (5), where $F\in\mathbb{R}^{n\times n}$ is positive real and $B\in\mathbb{R}^{m\times n}$ is full row rank, let $H_{1}$ and $N$ be the symmetric and skew-symmetric parts, respectively, of $F$, as in (8). Let $H_{2}$ be a symmetric positive definite matrix, such that the three conditions in (9)–(10) are satisfied. Finally, let $H$ to be the block-diagonal matrix defined in (7). Then, for $\alpha$ sufficiently small, GMRES with the block-diagonal preconditioner $M_{D}$ under the $H$-norm for left preconditioning or $H^{-1}$-norm for right preconditioning will converge in a fixed number of iterations independently of the dimensions of $K$.
@@ 3.3 Inexact Preconditioning
$\|\tilde{M}_{U}^{-1}K\|_{H}\leq\|\tilde{M}_{U}^{-1}M_{U}\|_{H}\|M_{U}^{-1}K\|_{H}\leq(1+C_{3}{\alpha})\|M_{U}^{-1}K\|_{H}\lesssim 1.$
@@
$\lesssim\alpha+2\|P_{1}^{-1}F-I\|_{H_{1}}\|M_{U}^{-1}K\|_{H}$
$\lesssim\alpha.$
Thus, if ${\alpha}$ is small enough, condition (2d) is satisfied and the iterative solver with $\tilde{M}_{U}$ as a preconditioner will converge in a fixed number of iterations.
@@ 4.1 Navier-Stokes Equations
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain. The Navier-Stokes equations with pure Dirichlet boundary conditions are [12]
$-\nu\Delta\mathbf{u}+(\mathbf{u}\cdot\nabla)\mathbf{u}+\nabla p=f\quad\text{in }\Omega,$
[…]
@@
where $\nu$ is a viscosity coefficient, $\mathbf{u}$ is the velocity, $p$ is the pressure, and $\mathbf{u}=\mathbf{u}_{d}$ provides the Dirichlet boundary conditions.
@@
$-\nu\Delta\mathbf{u}+(\mathbf{b}\cdot\nabla)\mathbf{u}+\nabla p=f\quad\text{in }\Omega,$
[…]
@@
$a(\mathbf{u},\mathbf{v})+b(\mathbf{v},p)$ $={\mathbf{f}}(\mathbf{v})\quad\forall\mathbf{v}\in\mathbf{V},$
@@
and ${\mathbf{f}}$ is the linear functional $\int_{\Omega}\mathbf{f}\cdot\mathbf{v}.$
@@
$\begin{bmatrix}F&B^{T}\\ B&0\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ p\end{bmatrix}=\begin{bmatrix}\mathbf{f}\\ 0\end{bmatrix},$ (17)
@@
$\begin{bmatrix}H_{1}+\frac{1}{\nu}N&B^{T}\\ B&0\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \frac{1}{\nu}p\end{bmatrix}=\begin{bmatrix}\frac{1}{\nu}\mathbf{f}\\ 0\end{bmatrix}.$
@@
We numerically solve the regularized lid-driven cavity problem using IFISS [13] to illustrate our results. The domain $\Omega$ is chosen as $[-1,1]^{2}$. Zero boundary conditions are imposed, except we take $u_{x}=1-x^{4}$ on $\{y=1,-1\leq x\leq 1\}$.
We set $\nu=1$, because our analysis requires it to be relatively large, and apply the Picard iteration. Since $\nu$ is relatively large, the nonlinear iterations converge quickly; we record average iteration counts and examine the performance of the linear solvers. We note that we have observed no significant differences among the linear solver iteration counts throughout the nonlinear iteration. We use the diagonal preconditioner $M_{D}$ defined in (16) and the upper triangular preconditioner $M_{U}$ defined in (11). We use left preconditioning for both; the results for right preconditioning with $M_{L}$ defined in (12) are virtually the same.
@@
System Size $\ell_{2}$ norm $H$-norm
@@
Table 1: Average iteration counts for Navier-Stokes
For the diagonal preconditioner, we have computed the parameters of Lamma 2.4 and have observed that $b\approx 2.25$ and $c\approx 0.016$. For the upper-triangular preconditioner, $b\approx 2.06$ and $c\approx 0.035$. In both cases we have $bc<1$, as required.

### v2

@@ 3 Preconditioned Saddle-Point Matrices with Zero in the Field of Values
where $F\in\mathbb{R}^{n\times n}$ is nonsingular and $B\in\mathbb{R}^{m\times n}$ is full row rank. We assume that $F$ is nonsymmetric and positive real (or positive definite), namely, that $\mathbf{u}^{T}F\mathbf{u}>0$ for all $\mathbf{u}\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}$.
@@
$\|N\|_{H_{1},H_{1}^{-1}}\leq{\eta},$ (11)
where $\eta$} is a constant independent of the matrix dimensions.
@@
The result can be readily deduced by using standard properties of norms; see [24, Lemma 2.1]. We have
[…]
@@
$\qquad\|B\|_{H_{1},H_{2}^{-1}}\leq C_{1},\qquad\min_{\mathbf{x}}\frac{\|B^{T}\mathbf{x}\|_{H_{1}^{-1}}}{\|\mathbf{x}\|_{H_{2}}}\geq C_{2},$ (12)
where $C_{1}$ and $C_{2}$ are independent of $\eta$} and the dimensions of $B$.
@@
[…]
$\geq C_{2}^{2}\min_{\mathbf{v}\in\mathbb{R}^{m}\backslash\{\mathbf{0}\}}\frac{\mathbf{v}^{T}BF^{-1}B^{T}\mathbf{v}}{\mathbf{v}^{T}BH_{1}^{-1}B^{T}\mathbf{v}}.$
Using [24, Lemma 3.4] and Lemma 3.3, we have
[…]
@@
$\geq\frac{1}{(1+{\eta})^{2}},$
where $\lambda_{k}(\cdot)$ denotes the $k$th eigenvalue of the input matrix.
@@
Let $\mathcal{T}$ denote a set of matrices. For a given scalar $\tau>0$, we say that $\|T\|\lesssim\tau$ for $T\in\mathcal{T}$ if there exists some constant $C>0$ independent of $T$ such that $\|T\|\leq C\tau$ for all $T\in\mathcal{T}$.
Definition 3.8.
Consider saddle-point matrices $K$ defined in (7), where $F$ is nonsymmetric and positive real, $B$ has full rank, and $H_{1},H_{2}$ and $N$ are as defined in Eq. (9) and in Definition 3.1. We define $\mathcal{K}_{\eta,C_{1},C_{2}}$ as the set of all such saddle-point matrices that satisfy Assumptions 3.2 and 3.5 for given values of $\eta,C_{1},$ and $C_{2}$,
In the convergence analysis that follows, our results are for matrices $K$ in the $\mathcal{K}_{\eta,C_{1},C_{2}}$ set.
@@ 3.1 Block-Triangular Preconditioners
It is well known that there are some differences in the use of left and right preconditioners. For example, in flexible GMRES it is necessary to use right preconditioning. The correct norm considered in GMRES for finite element discretizations is typically $\|\cdot\|_{H^{-1}}$ [1].
@@
$(M_{U}^{-1}K)^{-1}=\begin{bmatrix}I-F^{-1}B^{T}S^{-1}B&F^{-1}B^{T}S^{-1}H_{2}\\ S^{-1}B&I-S^{-1}H_{2}\end{bmatrix}$,
where we recall from (8) that $S=BF^{-1}B^{T}$. We now need to establish (3) and (4a)–(4c) in Lemma 2.4.
@@
The proof follows from putting together the bounds of Lemmas 3.3, 3.4, and 3.6.
@@
$\leq(1+2C_{1}^{2}{(1+{\eta})^{2}/C_{2}^{2}})+({(1+{\eta})^{2}/C_{2}^{2}}C_{1})+1+{(1+{\eta})^{2}/C_{2}^{2}}$
$\lesssim 1,$
where in the transition from the second to the third inequality we have used
$\|I-F^{-1}B^{T}S^{-1}B\|_{H_{1}}\leq 1+\|F^{-1}\|_{{H_{1}^{-1},H_{1}}}\|B^{T}\|_{H_{2},H_{1}^{-1}}\|S^{-1}\|_{H_{2}^{-1},H_{2}}\|B\|_{H_{1},H_{2}^{-1}},$
and similarly for the other terms.
@@
When ${\eta}<\frac{1}{2}$, we have
$\left\|H(M_{U}^{-1}K)-(M_{U}^{-1}K)^{T}H\right\|_{H,H^{-1}}\lesssim{\eta}.$ (15)
@@
When ${\eta}<\frac{1}{2}$ we have
$\|B_{12}\|_{2}\leq C_{1}\frac{\|N\|_{H_{1},H_{1}^{-1}}}{1-\|N\|_{H_{1},H_{1}^{-1}}}\leq C_{1}{\eta}/(1-{\eta})\leq 2C_{1}{\eta}\lesssim{\eta}$ (17)
@@
$\leq 2C_{1}^{2}{\eta}$
$\lesssim{\eta}.$
Substituting the above inequalities into (16), we obtain (15), as required. ∎
In order to be able to state an upcoming convergence theorem in precise terms, it is useful to rephrase the results of Lemmas 3.93.11 using some specific constants. Suppose $K\in\mathcal{K}_{\eta,C_{1},C_{2}}$ for some particular values of ${\eta}<\frac{1}{2}$, $C_{1}$, and $C_{2}$. Then Lemmas 3.93.11 say that there exist constants $C_{a}^{U},C_{b}^{U},C_{c}^{U}>0$ independent of the matrix dimensions such that
$\|M_{U}^{-1}K\|_{H}\leq C_{a}^{U},\quad\|(M_{U}^{-1}K)^{-1}\|_{H}\leq C_{b}^{U},$ (18)
$\left\|H(M_{U}^{-1}K)-(M_{U}^{-1}K)^{T}H\right\|_{H,H^{-1}}\leq C_{c}^{U}\,{\eta}.$
@@
Given a tolerance $\epsilon>0$ and a saddle-point system with matrix $K\in\mathcal{K}_{\eta,C_{1},C_{2}}$ for some particular values of ${\eta}<\frac{1}{2},C_{1}$, and $C_{2}$, let $H_{1}$ and $N$ be the symmetric and skew-symmetric parts, respectively, of $F$, as in (10). Let $H_{2}$ be a symmetric positive definite matrix. Finally, let $H$ be the block-diagonal matrix defined in (9). Then, using the constants defined in (18), if $C_{b}^{U}C_{c}^{U}\,{\eta}<1$ (i.e., (3) holds), then for any system that satisfies these requirements with $\eta$}, $C_{1}$, $C_{2}$, $C_{a}^{U}$, $C_{b}^{U}$, and $C_{c}^{U}$, there exists some integer $m\geq 1$ that depends on their values and on $\epsilon$, such that preconditioned GMRES with $M_{U}$ as a left preconditioner will converge (in the $H$-norm) to $\epsilon$ in no more than $m$ iterations.
@@
Trivially, by Lemmas 3.10 and 3.11, (3) holds when $\eta$} is sufficiently small. Lemmas 3.93.11 validate conditions (4a)–(4c). ∎
We now consider the right preconditioner $M_{L}$ defined in (14). The analysis is very similar to the left preconditioner case. We first define three inequalities analogous to (18):
$\|KM_{L}^{-1}\|_{H^{-1}}\leq C_{a}^{L},\quad\|(KM_{L}^{-1})^{-1}\|_{H^{-1}}\leq C_{b}^{L},$ (19)
$\left\|H^{-1}(KM_{L}^{-1})-(KM_{L}^{-1})^{T}H^{-1}\right\|_{H^{-1},H}\leq C_{c}^{L}\,{\eta}.$
We now present a theorem analogous to Theorem 3.12.
@@
Given a tolerance $\epsilon>0$ and a saddle-point system with matrix $K\in\mathcal{K}_{\eta,C_{1},C_{2}}$ for some particular values of ${\eta}<\frac{1}{2},C_{1}$, and $C_{2}$, let $H_{1}$ and $N$ be the symmetric and skew-symmetric parts, respectively, of $F$, as in (10). Let $H_{2}$ be a symmetric positive definite matrix. Let $H$ be the block-diagonal matrix defined in (9). Finally, assume we have the constants and inequalities defined in (19). Then, if $C_{b}^{L}C_{c}^{L}\,{\eta}<1$ (i.e., (3) holds), then for any system that satisfies these requirements with $\eta$}, $C_{1}$, $C_{2}$, $C_{a}^{L}$, $C_{b}^{L}$, and $C_{c}^{L}$, there exists some integer $m\geq 1$ that depends on their values and on $\epsilon$, such that preconditioned GMRES with $M_{L}$ as a right preconditioner will converge (in the $H^{-1}$-norm) to $\epsilon$ in no more than $m$ iterations.
@@ 3.2 A Block-Diagonal Preconditioner
is interesting in the context of this work, because contrary to block-triangular preconditioners, where one might select either an upper block-triangular preconditioner or a lower block-triangular preconditioner along with right or left preconditioning to avoid a situation of having zero in the field of values, here it is immediate that the field of values contains zero regardless of any such choices made. There is no practical difference between left and right preconditioning here, and we proceed with left preconditioning below. The preconditioned matrix is
@@
The analysis is essentially identical to that of Section 3.1.
@@
When ${\eta}<\frac{1}{2}$, we have
$\left\|H({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M_{D}}^{-1}K)-({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M_{D}}^{-1}K)^{T}H\right\|_{H,H^{-1}}\lesssim{\eta}.$
@@
Here we need two sets of inequalities, analogous both to (18) and (19):
$\|M_{D}^{-1}K\|_{H}\leq C_{a}^{D_{1}},\quad\|(M_{D}^{-1}K)^{-1}\|_{H}\leq C_{b}^{D_{1}},$ (21)
$\left\|H(M_{D}^{-1}K)-(M_{D}^{-1}K)^{T}H\right\|_{H,H^{-1}}\leq C_{c}^{D_{1}}\,{\eta}$
and
$\|KM_{D}^{-1}\|_{H^{-1}}\leq C_{a}^{D_{2}},\quad\|(KM_{D}^{-1})^{-1}\|_{H^{-1}}\leq C_{b}^{D_{2}},$ (22)
$\left\|H^{-1}(KM_{D}^{-1})-(KM_{D}^{-1})^{T}H^{-1}\right\|_{H^{-1},H}\leq C_{c}^{D_{2}}\,{\eta}.$
The convergence theorem is then given as follows.
@@
Given a tolerance $\epsilon>0$ and a saddle-point system with matrix $K\in\mathcal{K}_{\eta,C_{1},C_{2}}$ for some particular values of ${\eta}<\frac{1}{2},C_{1}$, and $C_{2}$, let $H_{1}$ and $N$ be the symmetric and skew-symmetric parts, respectively, of $F$, as in (10). Let $H_{2}$ be a symmetric positive definite matrix. Let $H$ be the block-diagonal matrix defined in (9). Finally, assume we have the constants and inequalities defined in (21) and (22). Then, if $C_{b}^{D_{1}}C_{c}^{D_{1}}\,{\eta}<1$ $C_{b}^{D_{2}}C_{c}^{D_{2}}\,{\eta}<1$ (i.e., (3) holds for both), then for any system that satisfies these requirements with $\eta$}, $C_{1}$, $C_{2}$, $C_{a}^{D_{1}}$, $C_{a}^{D_{2}}$, $C_{b}^{D_{1}}$, $C_{b}^{D_{2}}$, $C_{c}^{D_{1}}$, and $C_{c}^{D_{2}}$ there exists some integer $m\geq 1$ that depends on their values and on $\epsilon$ such that preconditioned GMRES with $M_{D}$ as a right preconditioner will converge in the $H^{-1}$-norm and preconditioned GMRES with $M_{D}$ as a left preconditioner will converge in the $H$-norm to $\epsilon$ in no more than $m$ iterations.
@@ 3.3 Inexact Preconditioning
$\|\tilde{M}_{U}^{-1}K\|_{H}\leq\|\tilde{M}_{U}^{-1}M_{U}\|_{H}\|M_{U}^{-1}K\|_{H}\leq(1+C_{3}{\eta})\|M_{U}^{-1}K\|_{H}\lesssim 1.$
@@
$\lesssim{\eta}+2\|P_{1}^{-1}F-I\|_{H_{1}}\|M_{U}^{-1}K\|_{H}$
$\lesssim{\eta}.$
Thus, if $\eta$} is small enough, (3) is satisfied and the iterative solver with $\tilde{M}_{U}$ as a preconditioner will converge in a fixed number of iterations.
@@ 4.1 Navier-Stokes Equations
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain. The Navier-Stokes equations with pure Dirichlet boundary conditions are
$-\nu\Delta\mathbf{u}+(\mathbf{u}\cdot\nabla)\mathbf{u}+\nabla p=\mathbf{f}\quad\text{in }\Omega,$
$\nabla\cdot\mathbf{u}=0\quad\text{in }\Omega,$
@@
where $\nu$ is a viscosity coefficient, $\mathbf{u}$ is the velocity, $p$ is the pressure, and $\mathbf{u}=\mathbf{u}_{d}$ provides the Dirichlet boundary conditions. For an in-depth description of the finite element solution of the Navier-Stokes equations, see, for example, [14, 16].
@@
$-\nu\Delta\mathbf{u}+(\mathbf{b}\cdot\nabla)\mathbf{u}+\nabla p=\mathbf{f}\quad\text{in }\Omega,$
$\nabla\cdot\mathbf{u}=0\quad\text{in }\Omega,$
@@
$a(\mathbf{u},\mathbf{v})+b(\mathbf{v},p)$ $=f(\mathbf{v})\quad\forall\mathbf{v}\in\mathbf{V},$
@@
and $f({\mathbf{v}})=\int_{\Omega}\mathbf{f}\cdot\mathbf{v}.$
@@
$\begin{bmatrix}F&B^{T}\\ B&0\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \mathbf{p}\end{bmatrix}=\begin{bmatrix}\mathbf{f}\\ \mathbf{0}\end{bmatrix}$, (23)
@@
$\begin{bmatrix}H_{1}+\frac{1}{\nu}N&B^{T}\\ B&0\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ \frac{1}{\nu}\mathbf{p}\end{bmatrix}=\begin{bmatrix}\frac{1}{\nu}\mathbf{f}\\ \mathbf{0}\end{bmatrix}$.
@@
Remark 4.1.
The skew-symmetric part is small in norm by Assumption 3.2, which holds if $\nu$ is large, but the saddle-point matrix is nonnormal and performing convergence analysis for GMRES is challenging. An analysis based on eigenvalue and eigenvector conditioning may be possible, based on observations and insights such as those in [15, Corollary 2.2 and Section 4.1] or [17], although we are not aware of a comprehensive analysis of this type for the Navier-Stokes equations. The field-of-values analysis is an alternative approach, applicable under the assumptions made.
Our setting can be extended to mixed boundary conditions. The bilinear form can be writen as
$a_{\rm mix}(\mathbf{u},\mathbf{v})=\nu\int_{\Omega}\nabla\mathbf{u}\cdot\nabla\mathbf{v}+\frac{1}{2}\int_{\Gamma_{n}}(\mathbf{b}\cdot\mathbf{n})\mathbf{u}\cdot\mathbf{v}+\int_{\Omega}(\mathbf{b}\cdot\nabla\mathbf{u})\cdot\mathbf{v}-\frac{1}{2}\int_{\Gamma_{n}}(\mathbf{b}\cdot\mathbf{n})\mathbf{u}\cdot\mathbf{v}.$
Here, $\Gamma_{n}$ is the boundary with Robin conditions $\mathbf{u}\cdot\mathbf{n}=g$ and $\mathbf{n}$ is a vector tangent to the boundary. Let us denote the leading block in the saddle-point system in this case by $\tilde{F}=\nu\tilde{H}_{1}+\tilde{N}$. Then
$a_{\rm mix}(\mathbf{u}_{h},\mathbf{v}_{h})=(\tilde{F}\mathbf{u}_{1},\mathbf{v}_{1})=\nu(\tilde{H}_{1}\mathbf{u}_{1},\mathbf{v}_{1})+(\tilde{N}\mathbf{u}_{1},\mathbf{v}_{1}),$
where
$(\tilde{H}_{1}\mathbf{u}_{1},\mathbf{v}_{1})=\int_{\Omega}\nabla\mathbf{u}_{1}\cdot\nabla\mathbf{v}_{1}+\frac{1}{2\nu}\int_{\Gamma_{n}}(\mathbf{b}\cdot\mathbf{n})\mathbf{u}_{1}\cdot\mathbf{v}_{1}$
and
$\int_{\Omega}(\mathbf{b}\cdot\nabla\mathbf{u}_{1})\cdot\mathbf{v}_{1}-\frac{1}{2}\int_{\Gamma_{n}}(\mathbf{b}\cdot\mathbf{n})\mathbf{u}_{1}\cdot\mathbf{v}_{1}=(\tilde{N}\mathbf{u}_{1},\mathbf{v}_{1}).$
Note that if $\nu$ is large enough, $\tilde{H}_{1}$ is spectrally equivalent to $H_{1}$, thus the mixed boundary conditions give us similar results to the setting with Dirichlet boundary conditions.
We numerically solve the regularized lid-driven cavity problem using IFISS [13] to illustrate our results. The domain $\Omega$ is chosen as $1,1]^{2}$. Zero boundary conditions are imposed, except we take $u_{x}=1-x^{4}$ on $\{y=1,-1\leq x\leq 1\}$.
We set $\nu=1$, because our analysis requires it to be relatively large, and apply the Picard iteration, using the IFISS default nonlinear tolerance. Since $\nu$ is relatively large, the nonlinear iterations converge quickly; we record average iteration counts and examine the performance of the linear solvers. We note that we have observed no significant differences among the linear solver iteration counts throughout the nonlinear iteration. We use the diagonal preconditioner $M_{D}$ defined in (20) and the upper triangular preconditioner $M_{U}$ defined in (13). We use left preconditioning for both; the results for right preconditioning with $M_{L}$ defined in (14) are virtually the same.
@@
System Size $\ell_{2}$-norm $H$-norm
@@
Table 1: Average iteration counts for Navier-Stokes. For each Picard iteration, the inner linear system solve was terminated when a relative residual tolerance of $\frac{\|r_{k}\|}{\|r_{0}\|}<10^{-5}$ was reached, where the norms used were the ones corresponding to results reported in the table: $\ell_{2}$-norm on the left and $H$-norm on the right.
For the diagonal preconditioner, we have computed the parameters of Lemma 2.4 in the $H$-norm and have observed that $b\approx 2.25$ and $c\approx 0.016$. For the upper-triangular preconditioner, $b\approx 2.06$ and $c\approx 0.035$. In both cases we have $bc<1$, as required.